### v1

@@ Acknowledgements
We thank present and past Prakash lab members for their helpful suggestions and support. We thank Zak Swartz (Marine Biological laboratory) and Athula Wikramanayake (University of Miami) for help with sea star culturing. We thank Phillip Gilette (University of Miami Rosenstiel School) for providing the sea urchin larvae. We thank Nitto Inc. (Atlanta, GA) for generously providing the spacer tapes used in our experiments. We thank Maciej Lisicki (University of Warsaw) and Lyndon Koens (University of Hull) for their invaluable suggestions on the theoretical modeling. V.N.P. thanks Manu Prakash (Stanford University) and William Gilpin (University of Texas at Austin) for many insightful discussions over the years.
@@
Data availability The data that support the results from this study are available from the corresponding author on request.
Code availability The computer codes that support the results from this study are available from the corresponding author on request.

### v2

@@ Acknowledgements
We thank present and past Prakash lab members for their helpful suggestions and support. We thank Zak Swartz (Marine Biological laboratory) and Athula Wikramanayake, Nat Clarke, Christopher Roden (University of Miami) for help with sea star culturing and other helpful suggestions. We thank Phillip Gilette (University of Miami Rosenstiel School) for providing the sea urchin larvae. We thank Nitto Inc. (Atlanta, GA) for generously providing the double-sided spacer tapes used in our experiments. We thank Maciej Lisicki (University of Warsaw) and Lyndon Koens (University of Hull) for their invaluable suggestions on the theoretical modeling. V.N.P. thanks Manu Prakash (Stanford University) and William Gilpin (University of Texas at Austin) for many insightful discussions over the years.
@@
Data availability Datasets associated with the manuscript will be deposited to Zenodo (link to be provided after completion of peer review).
Code availability The computer codes that support the results from this study will be deposited to Zenodo (link to be provided after completion of peer review).